OpenStudy (anonymous):

Could someone show me how to expand (2x + 3/y)^4 with the steps, please? @MIT 18.02 Multiva…

OpenStudy (anonymous):

Expand using the binomial theorem:\[(2x+\frac{3}{y})^4=\left(\begin{matrix}4 \\ 0\end{matrix}\right)(2x)^4(\frac{3}{y})^0+\left(\begin{matrix}4 \\ 1\end{matrix}\right)(2x)^3(\frac{3}{y})^1+\left(\begin{matrix}4 \\ 2\end{matrix}\right)(2x)^2(\frac{3}{y})^2\]\[+\left(\begin{matrix}4 \\ 3\end{matrix}\right)(2x)^1(\frac{3}{y})^3+\left(\begin{matrix}4 \\ 4\end{matrix}\right)(2x)^0(\frac{3}{y})^4\]

OpenStudy (anonymous):

Ok, just a quick question. For like (3/y)^2, how would i distribute the powers? Would it result in something like 3^2/y^2 ?

myininaya (myininaya):